# Equivalence of entropy solutions and gradient flows for pressureless 1D Euler systems

### v2

@@ 4.2.1. The Rankine–Hugoniot- and Oleĭnik E-condition
For piecewise smooth solutions of (4.4), the requirements for weak and entropy solutions can be reduced to conditions on the discontinuities, cf. [25, Chapter 2]. To this end, assume $M(t,x)$ takes the values $M_{l}(t)$ and $M_{r}(t)$ to the left and right of some shock curve $\Gamma=\{(t,x)\colon x=s(t)\}$. Denoting by $[[M]]$ the jump in $M$ across the curve $\Gamma$, and correspondingly $[[\eta(M)]]$, $[[q(t,M)]]$ for the other quantities, this leads to the inequality
@@ 4.2.2. Existence and uniqueness results
On the other hand, in cases such as Examples 4.3 and 4.4 the part involving the forcing term $f_{\rho}$ simplifies, and we we are led to consider flux functions of the form
@@
A function $M\colon\mathbb{R}\to\mathbb{R}$ belongs to the set $\mathscr{R}$ if it is nondecreasing, right-continuous and has limits 00 and $1$ at respectively $\infty$- ∞ and +$\infty$+ ∞. Moreover, $M\in\mathscr{R}$ belongs to $\mathscr{R}_{p}$ for $p\geq 1$ if its distributional derivative $\partial_{x}M$ has finite $p$-moment, that is, $\int_{\mathbb{R}}\lvert x\rvert^{p}\partial_{x}M$ is finite.
@@
The argument follows the lines of the proof of [25, Theorem 7.10] or [19, Theorem 4.3.1], with some adjustments to cater for our non-Lipschitzian flux function. Let $\chi^{\varepsilon}(t)$ for $t\in[0,T]$ be a smooth approximation of the characteristic function for the interval $[s,t]\subset[0,T]$ such that $\lim\limits_{\varepsilon\to 0+}\chi^{\varepsilon}=\chi_{[s,t]}$. Furthermore we choose $\varphi_{\varepsilon}(t,x)=\chi^{\varepsilon}(t)\phi(x)$ for an arbitrary $\phi\in C_{\mathrm{c}}^{\infty}(\mathbb{R})$. For this test function, the weak formulation of (4.4) reads
@@
$-\int_{\mathbb{R}}\phi(x)(M(t,x)-M(s,x))\mathop{}\!\mathrm{d}x+\int_{s}^{t}\int_{\mathbb{R}}\phi^{\prime}(x)\mathcal{U}(r,M(r,x))\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}r=0.$
@@
$\leq\sum_{i=0}^{n}\lvert B_{i}(t)\rvert\left(\int_{\mathbb{R}}\lvert\phi(x)\rvert^{p^{\prime}}\mathop{}\!\mathrm{d}\rho(t,x)\right)^{\frac{1}{p^{\prime}}}\lVert\mathcal{F}_{i,M}^{\prime}\rVert_{L^{p}(\mathbb{R},\rho)}\leq\sum_{i=0}^{n}\lvert B_{i}(t)\rvert\lVert F_{i}\rVert_{L^{p}(0,1)}.$
Taking supremum over $\lvert\phi\rvert\leq 1$ and combining with the previous estimate we obtain (4.22). Note that we could also have arrived at the same bound by estimating the total variation of $\mathcal{U}(t,M(t,\cdot))$ as in the proof of Lemma 4.2. Since $\mathscr{R}\subset BV(\mathbb{R})$, then as in the proof of [19, Theorem 4.3.1] it follows from the characterization of $BV_{\mathrm{loc}}$-functions in [19, Theorem 1.7.2] and (4.22) that $M\in BV_{\mathrm{loc}}([0,T]\times\mathbb{R})$. ∎
@@
$-\int_{\mathbb{R}}\phi^{\prime}(x)(M(t,x)-M(s,x))\mathop{}\!\mathrm{d}x+\int_{s}^{t}\int_{\mathbb{R}}\phi^{\prime\prime}(x)\mathcal{U}(r,M(r,x))\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}r=0,$
and integrating by parts we arrive at
@@
having applied Young’s inequality. Now, define $\phi_{R}(x)\coloneqq(\lvert x\rvert\wedge R)^{p}$ such that $\phi^{\prime}_{R}(x)=p\lvert x\rvert^{p-1}\operatorname*{sgn}(x)$ for $\lvert x\rvert<R$ and $\phi^{\prime}_{R}(x)=0$ for $\lvert x\rvert>R$. Since $\phi^{\prime}_{R}$ is compactly supported we can find a sequence of smooth approximations $(\phi^{\varepsilon}_{R})^{\prime}$ such that $(\phi^{\varepsilon}_{R})^{\prime}(x)=\phi^{\prime}_{R}(x)$ for $\lvert x\rvert\leq R-\varepsilon$, $(\phi^{\varepsilon}_{R})^{\prime}(x)=0$ for $\lvert x\rvert\geq R$, and $(\phi^{\varepsilon}_{R})^{\prime}(x)$ tends to $\phi_{R}^{\prime}(x)$ from below (above) for $R-\varepsilon<x<R$ ($-R<x<\varepsilon-R$). In consequence, the primitive $\phi^{\varepsilon}_{R}(x)=\int_{-\infty}^{x}(\phi^{\varepsilon}_{R})^{\prime}(y)\mathop{}\!\mathrm{d}y$ tends to $\phi_{R}(x)-R^{p}\in C_{\mathrm{c}}(\mathbb{R})$ pointwise from below. By the dominated convergence theorem it then follows that
@@ 4.3. Entropy solutions of the Euler system
Let $p\geq 2$ and assume that $M\in C([0,T];L^{1}_{\mathrm{loc}}(\mathbb{R}))$, $M\in\mathscr{R}$ for $t\in(0,T)$ and that it satisfies (4.4) in distributions $\mathscr{D}^{\prime}((0,T)\times\mathbb{R})$ with $\mathcal{U}(t,m)$ given by (4.18). According to Proposition 4.11 we let
$\mathcal{U}^{\prime}_{M}\partial_{x}M(t,\cdot)=\partial_{x}\mathcal{U}(t,M)\enspace\text{ for every }t\in(0,T).$
Then we have $M$, $\mathcal{U}(t,M)\in BV_{\mathrm{loc}}((0,T)\times\mathbb{R})$, $\mathcal{U}^{\prime}_{M}\in L^{1}(\lvert\nabla M\rvert)$ and
@@
where $\nabla=(\partial_{t},\partial_{x})$ denotes the $BV$-gradient.
@@
Let $p\in[2,\infty)$, $(\rho_{0},v_{0})\in\mathscr{T}_{p}(\mathbb{R})$ and denote by $M^{0}$ the right-continuous cumulative distribution function of $\rho_{0}$. Let $M\in\mathscr{R}_{p}$ be the corresponding entropy solution of (4.4) with flux function (4.18). Define $\rho(t,x)\coloneqq\partial_{x}M(t,x)$ and $v(t,x)\coloneqq\mathcal{U}_{M}^{\prime}(t,x)$ as in Theorem 4.13. Then $(\rho(t,\cdot),v(t,\cdot))\in\mathscr{T}_{p}(\mathbb{R})$ for $t\in(0,T)$ and $(\rho,v)$ is a distributional solution of (1.1) for
@@
On the other hand, if $\alpha<0$ then the flux function is strictly convex on $[M_{l},M_{r}]$ for all time, meaning $\mathcal{U}_{\smallsmile}(t,m)=\mathcal{U}(t,m)$. For fixed $t$ we consider the derivative of $m\mapsto\mathcal{U}_{\smallsmile}(t,m)$, which we denote by $\mathcal{U}^{\prime}_{\smallsmile}(t,m)$. Since $m\mapsto\mathcal{U}^{\prime}_{\smallsmile}(t,m)$ is monotone in $m$ we may define its inverse $v\mapsto(\mathcal{U}^{\prime}_{\smallsmile})^{-1}(t,v)$. The left and right constant states will travel with respective velocities $\mathcal{U}^{\prime}(t,M_{l})=v^{0}-t\frac{\alpha}{2}(2M_{l}-1)$ and $\mathcal{U}^{\prime}(t,M_{r})>\mathcal{U}^{\prime}(t,M_{l})$. The remaining middle section will be a continuous transition from $m=M_{l}$ to $m=M_{r}$, where the velocity corresponding to $m$ is $\mathcal{U}^{\prime}(t,m)$. Integrating we find $x=v^{0}t-\frac{1}{4}t^{2}\alpha(2m-1)$, which we can invert to find the value of $M(t,x)$ in the middle section, and we find the solution
@@ 5.1. Projections and shock admissibility
where $[[M(x)]]=M(x)-M(x-)$. As in Section 3.2, we can consider the open set $\Omega_{X}$ as consisting of countably many disjoint maximal intervals, and if $m\notin\Omega_{X}$, then $M(t,X(t,m))=m$. On the other hand, if $m\in(m_{l},m_{r})$ for a maximal interval $(m_{l},m_{r})\subset\Omega_{X}$, then by the right-continuity of $X$ and $M$ we have $M(X(m))=m_{r}$ and $M(X(m)-)=m_{l}$. Combining these properties with the expression for $\mathcal{W}^{\prime}_{M}$, we find exactly $\operatorname{P}_{\mathscr{H}_{X}}W$ defined by (3.11). ∎
@@
(5.5) $\eta_{k,M}^{\prime}(x)$ =$=\begin{cases}\operatorname*{sgn}(M(x)-k),&\mu(\{x\})=0,\\[5.69054pt] \displaystyle\frac{\lvert M(x)-k\rvert-\lvert M(x-)-k\rvert}{\mu(\{x\})},&\mu(\{x\})\neq 0\end{cases}$
@@
(5.6) $q_{k,M}^{\prime}(t,x)$ =$=\begin{cases}\operatorname*{sgn}(M(x)-k)U(t,M(x)),&\mu(\{x\})=0,\\[5.69054pt] \displaystyle\frac{[[\operatorname*{sgn}(M-k)\left(\mathcal{U}(t,M)-\mathcal{U}(t,k)\right)]](x)}{\mu(\{x\})},&\mu(\{x\})\neq 0,\end{cases}$
@@
where we as short-hand notation write $[[f(t,M)]](x)\coloneqq f(t,M(x))-f(t,M(x-))$ for a function $f\colon[0,T]\times\mathbb{R}\to\mathbb{R}$. Note that we can apply Lemma 5.1 to (4.13) to obtain the identities
@@
Assume $M$ is a weak solution of (4.4) with a flux function of the form $\mathcal{U}(t,m)=\int_{0}^{m}U(t,\omega)\mathop{}\!\mathrm{d}\omega$ for $U(t)\in L^{2}(\Omega)$. Then it is an entropy solution if and only if it satisfies the Oleĭnik E-condition (4.17) along its discontinuities, i.e., where $M(t,x)>M(t,x-)$.
@@
00 0 $\leq-\int_{0}^{T}\int_{\mathbb{R}}\varphi(t,x)\eta^{\prime}_{k,M}(t,x)\partial_{t}M(t,x)\mathop{}\!\mathrm{d}t-\int_{0}^{T}\int_{\mathbb{R}}\varphi(t,x)q^{\prime}_{k,M}(t,x)\partial_{x}M(t,x)\mathop{}\!\mathrm{d}t$
@@ 5.2. Equivalence theorem and proof
We must now show that $U(t)-\operatorname{P}_{\mathscr{H}_{X(t)}}U(t)\in N_{X(t)}\mathscr{K}$ for a.e. $t$. However, since $M$ is an entropy solution, by Lemma 5.4 it satisfies the Oleĭnik E-condition (4.17) at any discontinuity of $M(t,x)$ for a.e. $t$, and so by (5.2) and Lemma 5.2 we conclude that $U(t)-V(t)$ is contained in the normal cone $N_{X(t)}\mathscr{K}$ for a.e. $t$. Then it follows that $V(t)=\operatorname{P}_{T_{X(t)}\mathscr{K}}U(t)$ for a.e. $t$. Therefore $X(t)$ solves the differential inclusion (3.21a) and is almost everywhere differentiable again due to (4.24). Then, as argued in the proof of [7, Theorem 3.5], see also [10, Remark 3.9], at every point of differentiability we must have $\dot{X}(t,m)=V(t,m)=\operatorname{P}_{\mathscr{H}_{X(t)}}U(t)$. It remains to show that $\lim\limits_{t\to 0+}X(t)=X^{0}$ in $L^{2}(\Omega)$, but this follows from the identity (4.20) and the time-continuity (4.24) for $p=2$. ∎
@@
where $h\colon\mathbb{R}\to\mathbb{R}$ is a given Lipschitz function, and $f\colon\mathbb{R}\to\mathbb{R}$ is a smooth, convex function satisfying $\lim_{\lvert v\rvert\to+\infty}f(v)/\lvert v\rvert=+\infty$. In particular, for $f(v)=\frac{1}{2}v^{2}$ we recover a version of (1.1). In their study they make use of a generalization of the Hopf–Lax–Oleĭnik formula, which is similar to the sticky Lagrangian solutions in Section 3 in the sense that there is a representation formula, see also [46]. An interesting observation is that their uniqueness results require $h$ to be nondecreasing, which means that its primitive is concave. If one, to make sense of the above system in the case of measures with Dirac parts, replaced $h(\int_{-\infty}^{x}\rho\mathop{}\!\mathrm{d}y)$ with its Vol’pert average (4.7), the Lagrangian representation satisfying (3.16) would be $F[X](m)=h(m)$, independent of $X$ as in the Euler–Poisson case. Furthermore, since $h$ is nondecreasing, using Lemma 3.1 one can then verify that this functional is sticking in the sense of Definition 3.4, meaning the unique Lagrangian solutions are globally sticky as in Definition 3.8.
@@
Compared to our result, we merely argue that the Lagrangian solutions of the attractive Euler–Poisson system studied here are also entropy solutions in the sense of [26]. Indeed, the weak continuity of the kinetic energy follows from the right-continuity of $V$, cf. Lemma 3.6, which then carries over to $\lVert V\rVert^{2}_{L^{2}(\Omega)}$. It remains to verify (1.8), and to that end, as noted in [35], it is sufficient that the map $m\mapsto X(t,m)-tV(t,m)$ is nondecreasing. Using the fact that (1.3) for $\alpha\geq 0$ has globally sticky solutions given by the projection formulas (3.22) and (3.23), together with (3.13) we find $X(t)-tV(t)=\operatorname{P}_{\mathscr{H}_{X(t)}}(Y(t)-tU(t))$. Since $\operatorname{P}_{\mathscr{H}_{X}}$ preserves monotonicity, it is sufficient to show that $Y-tU$ is nondecreasing, and we easily compute $Y(t)-tU(t)=X^{0}(m)+\frac{1}{2}t^{2}\alpha(m-\frac{1}{2})$ which in our attractive case $\alpha>0$ clearly is increasing.

### v3

@@ 4.2.1. The Rankine–Hugoniot- and Oleĭnik E-condition
For piecewise smooth solutions of (4.4), the requirements for weak and entropy solutions can be reduced to conditions on the discontinuities, cf. [25, Chapter 2]. To this end, assume $M(t,x)$ takes the values $M_{l}(t)$ and $M_{r}(t)$ to the left and right of some shock curve $\Gamma=\{(t,x)\colon x=s(t)\}$. Denoting by $[[M]]$ the jump in $M$ across the curve $\Gamma$, and correspondingly $\eta(M)]]$, $[[q(t,M)]]$ for the other quantities, this leads to the inequality
@@ 4.2.2. Existence and uniqueness results
On the other hand, in cases such as Examples 4.3 and 4.4 the part involving the forcing term $f_{\rho}$ simplifies, and we are led to consider flux functions of the form
@@
A function $M\colon\mathbb{R}\to\mathbb{R}$ belongs to the set $\mathscr{R}$ if it is nondecreasing, right-continuous and has limits 0 and $1$ at respectively -$\infty$ and $\infty$. Moreover, $M\in\mathscr{R}$ belongs to $\mathscr{R}_{p}$ for $p\geq 1$ if its distributional derivative $\partial_{x}M$ has finite $p$-moment, that is, $\int_{\mathbb{R}}\lvert x\rvert^{p}\partial_{x}M$ is finite.
@@
The argument follows the lines of the proof of [25, Theorem 7.10] or [19, Theorem 4.3.1], with some adjustments to cater for our non-Lipschitzian flux function. Let $\chi^{\varepsilon}(t)$ for $t\in[0,T]$ be a smooth approximation of the characteristic function for the interval $s,t]\subset[0,T]$ such that $\lim\limits_{\varepsilon\to 0+}\chi^{\varepsilon}=\chi_{[s,t]}$. Furthermore we choose $\varphi_{\varepsilon}(t,x)=\chi^{\varepsilon}(t)\phi(x)$ for an arbitrary $\phi\in C_{\mathrm{c}}^{\infty}(\mathbb{R})$. For this test function, the weak formulation of (4.4) reads
@@
$\int_{\mathbb{R}}\phi(x)(M(t,x)-M(s,x))\mathop{}\!\mathrm{d}x+\int_{s}^{t}\int_{\mathbb{R}}\phi^{\prime}(x)\mathcal{U}(r,M(r,x))\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}r=0.$
@@
$\leq\sum_{i=0}^{n}\lvert B_{i}(t)\rvert\left(\int_{\mathbb{R}}\lvert\phi(x)\rvert^{p^{\prime}}\mathop{}\!\mathrm{d}\rho(t,x)\right)^{\frac{1}{p^{\prime}}}\lVert\mathcal{F}_{i,M}^{\prime}\rVert_{L^{p}(\mathbb{R},\rho)}\leq\sum_{i=0}^{n}\lvert B_{i}(t)\rvert\lVert F_{i}\rVert_{L^{p}(0,1)},$
where we for an arbitrary time $t$ integrate over space. Taking supremum over $\lvert\phi\rvert\leq 1$ and combining with the previous estimate we obtain (4.22). Note that we could also have arrived at the same bound by estimating the total variation of $\mathcal{U}(t,M(t,\cdot))$ as in the proof of Lemma 4.2. Since $\mathscr{R}\subset BV(\mathbb{R})$, then as in the proof of [19, Theorem 4.3.1] it follows from the characterization of $BV_{\mathrm{loc}}$-functions in [19, Theorem 1.7.2] and (4.22) that $M\in BV_{\mathrm{loc}}([0,T]\times\mathbb{R})$. ∎
@@
$\int_{\mathbb{R}}\phi^{\prime}(x)(M(t,x)-M(s,x))\mathop{}\!\mathrm{d}x+\int_{s}^{t}\int_{\mathbb{R}}\phi^{\prime\prime}(x)\mathcal{U}(r,M(r,x))\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}r=0,$
and integrating by parts, omitting $x$ in the differentials for brevity, we arrive at
@@
having applied Young’s inequality. Now, define $\phi_{R}(x)\coloneqq(\lvert x\rvert\wedge R)^{p}$ such that $\phi^{\prime}_{R}(x)=p\lvert x\rvert^{p-1}\operatorname*{sgn}(x)$ for $\lvert x\rvert<R$ and $\phi^{\prime}_{R}(x)=0$ for $\lvert x\rvert>R$. Since $\phi^{\prime}_{R}$ is compactly supported we can find a sequence of smooth approximations $(\phi^{\varepsilon}_{R})^{\prime}$ such that $(\phi^{\varepsilon}_{R})^{\prime}(x)=\phi^{\prime}_{R}(x)$ for $\lvert x\rvert\leq R-\varepsilon$, $(\phi^{\varepsilon}_{R})^{\prime}(x)=0$ for $\lvert x\rvert\geq R$, and $(\phi^{\varepsilon}_{R})^{\prime}(x)$ tends to $\phi_{R}^{\prime}(x)$ from below (above) for $R-\varepsilon<x<R$ ($R<x<\varepsilon-R$). In consequence, the primitive $\phi^{\varepsilon}_{R}(x)=\int_{-\infty}^{x}(\phi^{\varepsilon}_{R})^{\prime}(y)\mathop{}\!\mathrm{d}y$ tends to $\phi_{R}(x)-R^{p}\in C_{\mathrm{c}}(\mathbb{R})$ pointwise from below. By the dominated convergence theorem it then follows that
@@ 4.3. Entropy solutions of the Euler system
Let $p\geq 2$ and assume that $\in C([0,T];L^{1}_{\mathrm{loc}}(\mathbb{R})), MM\in\mathscr{R}$ for $t\in(0,T)$ and that it satisfies (4.4) in distributions $\mathscr{D}^{\prime}((0,T)\times\mathbb{R})$ with $\mathcal{U}(t,m)$ given by (4.18). According to Proposition 4.11 we let
$\mathcal{U}^{\prime}_{M}\partial_{x}M(t,\cdot)=\partial_{x}\mathcal{U}(t,M)\kern 5.0pt\text{ for every }t\in(0,T).$
@@
(4.23) $\nabla\mathcal{U}(t,M)=\mathcal{U}_{M}^{\prime}\nabla M+\left(\sum_{i=1}^{n}b_{i}(t)\mathcal{F}_{i}(M),\>0\right)=\left(-(\mathcal{U}_{M}^{\prime})^{2},\>\mathcal{U}_{M}^{\prime}\right)\partial_{x}M+\left(\sum_{i=1}^{n}b_{i}(t)\mathcal{F}_{i}(M),\>0\right),$
where $\nabla=(\partial_{t},\partial_{x})$ denotes the BVBV-gradient.
@@
Let $p\in[2,\infty)$, $(\rho_{0},v_{0})\in\mathscr{T}_{p}(\mathbb{R})$ and denote by $M^{0}$ the right-continuous distribution function of $\rho_{0}$. Let $M\in\mathscr{R}_{p}$ be the corresponding entropy solution of (4.4) with flux function (4.18). Define $\rho(t,x)\coloneqq\partial_{x}M(t,x)$ and $v(t,x)\coloneqq\mathcal{U}_{M}^{\prime}(t,x)$ as in Theorem 4.13. Then $(\rho(t,\cdot),v(t,\cdot))\in\mathscr{T}_{p}(\mathbb{R})$ for $t\in(0,T)$ and $(\rho,v)$ is a distributional solution of (1.1) for
@@
On the other hand, if $\alpha<0$ then the flux function is strictly convex on $M_{l},M_{r}]$ for all time, meaning $\mathcal{U}_{\smallsmile}(t,m)=\mathcal{U}(t,m)$. For fixed $t$ we consider the derivative of $m\mapsto\mathcal{U}_{\smallsmile}(t,m)$, which we denote by $\mathcal{U}^{\prime}_{\smallsmile}(t,m)$. Since $m\mapsto\mathcal{U}^{\prime}_{\smallsmile}(t,m)$ is monotone in $m$ we may define its inverse $v\mapsto(\mathcal{U}^{\prime}_{\smallsmile})^{-1}(t,v)$. The left and right constant states will travel with respective velocities $\mathcal{U}^{\prime}(t,M_{l})=v^{0}-t\frac{\alpha}{2}(2M_{l}-1)$ and $\mathcal{U}^{\prime}(t,M_{r})>\mathcal{U}^{\prime}(t,M_{l})$. The remaining middle section will be a continuous transition from $m=M_{l}$ to $m=M_{r}$, where the velocity corresponding to $m$ is $\mathcal{U}^{\prime}(t,m)$. Integrating we find $x=v^{0}t-\frac{1}{4}t^{2}\alpha(2m-1)$, which we can invert to find the value of $M(t,x)$ in the middle section, and we find the solution
@@ 5.1. Projections and shock admissibility
where [[M(x)]]=M(x)M(x)[[M(x)]]=M(x)-M(x-). As in Section 3.2, we can consider the open set $\Omega_{X}$ as consisting of countably many disjoint maximal intervals, and if $m\notin\Omega_{X}$, then $M(t,X(t,m))=m$. On the other hand, if $m\in(m_{l},m_{r})$ for a maximal interval $(m_{l},m_{r})\subset\Omega_{X}$, then by the right-continuity of $X$ and $M$ we have $M(X(m))=m_{r}$ and $M(X(m)-)=m_{l}$. Combining these properties with the expression for $\mathcal{W}^{\prime}_{M}$, we find exactly $\operatorname{P}_{\mathscr{H}_{X}}W$ defined by (3.11). ∎
@@
[…]
@@
[…]
where we as short-hand notation write $f(t,M)]](x)\coloneqq f(t,M(x))-f(t,M(x-))$ for a function $f\colon[0,T]\times\mathbb{R}\to\mathbb{R}$. Note that we can apply Lemma 5.1 to (4.13) to obtain the identities
@@
Assume $M$ is a weak solution of (4.4) with a flux function of the form $\mathcal{U}(t,m)=\int_{0}^{m}U(t,\omega)\mathop{}\!\mathrm{d}\omega$ for $U(t)\in L^{2}(\Omega)$. Then it is an entropy solution if and only if it satisfies the Oleĭnik E-condition (4.17) along its discontinuities, i.e., where M(t,x)>M(t,x)M(t,x)>M(t,x-).
@@
$0$ $\leq-\int_{0}^{T}\int_{\mathbb{R}}\varphi(t,x)\eta^{\prime}_{k,M}(t,x)\partial_{t}M(t,x)\mathop{}\!\mathrm{d}t-\int_{0}^{T}\int_{\mathbb{R}}\varphi(t,x)q^{\prime}_{k,M}(t,x)\partial_{x}M(t,x)\mathop{}\!\mathrm{d}t$
@@ 5.2. Equivalence theorem and proof
We must now show that $U(t)-\operatorname{P}_{\mathscr{H}_{X(t)}}U(t)\in N_{X(t)}\mathscr{K}$ for a.e. $t$. However, since $M$ is an entropy solution, by Lemma 5.4 it satisfies the Oleĭnik E-condition (4.17) at any discontinuity of $M(t,x)$ for a.e. $t$, and so by (5.2) and Lemma 5.2 we conclude that U(t)V(t)U(t)-V(t) is contained in the normal cone $N_{X(t)}\mathscr{K}$ for a.e. $t$. Then it follows that $V(t)=\operatorname{P}_{T_{X(t)}\mathscr{K}}U(t)$ for a.e. $t$. Therefore $X(t)$ solves the differential inclusion (3.21a) and is almost everywhere differentiable again due to (4.24). Then, as argued in the proof of [7, Theorem 3.5], see also [10, Remark 3.9], at every point of differentiability we must have $\dot{X}(t,m)=V(t,m)=\operatorname{P}_{\mathscr{H}_{X(t)}}U(t)$. It remains to show that $\lim\limits_{t\to 0+}X(t)=X^{0}$ in $L^{2}(\Omega)$, but this follows from the identity (4.20) and the time-continuity (4.24) for $p=2$. ∎
@@
where $h\colon\mathbb{R}\to\mathbb{R}$ is a given Lipschitz function, and $f\colon\mathbb{R}\to\mathbb{R}$ is a smooth, convex function satisfying $\lim_{\lvert v\rvert\to+\infty}f(v)/\lvert v\rvert$$\infty$. In particular, for $f(v)=\frac{1}{2}v^{2}$ we recover a version of (1.1). In their study they make use of a generalization of the Hopf–Lax–Oleĭnik formula, which is similar to the sticky Lagrangian solutions in Section 3 in the sense that there is a representation formula, see also [46]. An interesting observation is that their uniqueness results require $h$ to be nondecreasing, which means that its primitive is concave. If one, to make sense of the above system in the case of measures with Dirac parts, replaced $h(\int_{-\infty}^{x}\rho\mathop{}\!\mathrm{d}y)$ with its Vol’pert average (4.7), the Lagrangian representation satisfying (3.16) would be $F[X](m)=h(m)$, independent of $X$ as in the Euler–Poisson case. Furthermore, since $h$ is nondecreasing, using Lemma 3.1 one can then verify that this functional is sticking in the sense of Definition 3.4, meaning the unique Lagrangian solutions are globally sticky as in Definition 3.8.
@@
Compared to our result, we merely argue that the Lagrangian solutions of the attractive Euler–Poisson system studied here are also entropy solutions in the sense of [26]. Indeed, the weak continuity of the kinetic energy follows from the right-continuity of $V$, cf. Lemma 3.6, which then carries over to $\lVert V\rVert^{2}_{L^{2}(\Omega)}$. It remains to verify (1.9), and to that end, as noted in [35], it is sufficient that the map $m\mapsto X(t,m)-tV(t,m)$ is nondecreasing. Using the fact that (1.4) for $\alpha\geq 0$ has globally sticky solutions given by the projection formulas (3.22) and (3.23), together with (3.13) we find $X(t)-tV(t)=\operatorname{P}_{\mathscr{H}_{X(t)}}(Y(t)-tU(t))$. Since $\operatorname{P}_{\mathscr{H}_{X}}$ preserves monotonicity, it is sufficient to show that YtUY-tU is nondecreasing, and we easily compute $Y(t)-tU(t)=X^{0}(m)+\frac{1}{2}t^{2}\alpha(m-\frac{1}{2})$ which in our attractive case $\alpha>0$ clearly is increasing.